Each of the normal students will demand courses in the amount:
QN= 100 – 0.25 P where Q is the number of course-hours per academic year and P is the price per course-hour. Each of the workaholics, on the other hand, has a demand of:
QW= 200 – 0.5 P The school will admit 180 students from each group. (Students may choose not to enroll.) The marginal cost to the School is constant at $100 per student course-hour. Sloan’s fixed cost is $2.0 million per year. For each of the following pricing schemes, please derive the optimal price structure and the total profit, on the assumption that the School’s objective is to maximize profit from the MBA program. a. A single fixed tuition per academic year (with a zero price per course-hour), with no limit on the number of course hours. Hint: Sloan may decide to induce both types of students to enroll or just the workaholics
1
Expert's answer
2018-04-09T02:53:11-0400
QN= 100 – 0.25 P, so P = 400 - 4QN,
QW= 200 – 0.5 P, so P = 400 - 2QW.
The school will admit 180 students from each group. MC = $100 per student course-hour. FC = $2.0 million per year.
Profit is maximized when MR = MC = P, so for normal students:
100 = 400 - 4QN,
QN = 75, P = $100.
For workaholics:
100 = 400 - 2QW,
QW = 150, P = $100.